The United States is China's largest export market. Imports from China grew from about 82 billion dollars in 1999 to 152 billion dollars in 2003. Let x=0 correspond to 1990, and estimate the amount of imports in 2006.
Answer by checkley77(12569)

(152-82)/(2003-1999)
70/4=17.5 BILLION IS THE GROWTH PER YEAR.
82+17.5=99.5 WOULD BE THE RATE FOR 2000.
(X-99.5)/6=17.5
X-99.5=6*17.5
X-99.5=105
X=105+99.5
X=204.5 BILLION WOULD BE THE IMPORTS FOR 2006.